# The Atacama Cosmology Telescope: DR6 Gravitational Lensing and SDSS BOSS cross-correlation measurement and constraints on gravity with the EG statistic

### v1

@@ I Introduction
While the current standard model of cosmology, CDM, fits a wide range of observational constraints (e.g. [1, 2, 3, 4, 5, 6]), the explanation behind the observed magnitude of the cosmological constant, , remains unresolved [7]. This has motivated the proposal of several alternative explanations, including a novel type of energy, called dark energy, and gravitational models including gravity [8], DGP [9], Chameleon gravity [10], TeVeS [11] and others (for a review see Clifton et al. [12]). These modified gravitational models are generally designed to match the observed homogeneous accelerated expansion of the late-time universe and employ screening mechanisms so that they match general relativity (GR) on solar system scales [13]. However, they do typically predict deviations from GR in the equations of motion affecting the growth of density fluctuations that seed large-scale structure (LSS) in the Universe [14]. Probing the growth history of the Universe, therefore, offers a powerful way to test gravity, allowing the construction and measurement of discriminating tests between CDM and alternative models of gravity (see e.g. [15, 13, 16, 17]).
@@
Photons from the Cosmic Microwave Background (CMB) get gravitationally lensed along the line of sight through the intervening gravitational potential formed from the large-scale structure of the Universe. This CMB lensing effect can be detected in the CMB anisotropies through its characteristic effect of coupling different scales which would otherwise be independent from each other. CMB lensing maps have been created for recent CMB measurements from the Planck satellite [18], the Atacama Cosmology Telescope (ACT) [19, 20] and the South Pole Telescope (SPT) [21, 22]. The newest ACT CMB lensing map constructed based on the Data Release 6 (DR6) CMB dataset covers a total of and offers significantly higher resolution and signal-to-noise than previous maps based on Planck data [20].
@@
The cosmic volume probed by CMB lensing along the line of sight overlaps with the volumes probed by galaxy surveys. This allows measurements of the correlation between CMB lensing and galaxy clustering which serves as an opportunity to break degeneracies between the amplitude of fluctuations and the galaxy bias for galaxy clustering alone. The Sloan Digital Sky Survey III Baryon Oscillation Spectroscopic Survey (SDSS BOSS) DR12 contains two large spectroscopic galaxy samples, CMASS and LOWZ which may be used to study this correlation [23, 24, 1]. The correlation has been previously measured with the smaller ACT DR4 CMB lensing map in Darwish et al. [19], Marques et al. [25]. Combining the SDSS BOSS galaxy catalogs with the ACT DR6 CMB lensing map offers the opportunity to derive new competitive constraints on the correlation between CMB lensing and galaxy clustering, independent of Planck, and allows new constraints on gravity. To demonstrate the accuracy of this correlation we perform a detailed suite of null and consistency tests to show that biases from systematic effects like extragalactic foregrounds and magnification bias are mitigated and accurately accounted for. For this work, we focus on using these cross-correlation measurements for constraints on gravity and make them available for further analysis.
@@
The measurement of this cross-correlation, combined with previous analyses of the galaxy data in Wenzl et al. [26], allow us to constrain the so-called statistic proposed in Zhang et al. [27]. This statistic builds a discriminating test of gravity by comparing the divergence of the peculiar velocity field with a measurement of gravitational lensing . The proposed ratio compares the two sides of the generalized Poisson equation and tests the equivalence of the spatial and temporal gravitation potentials, directly testing the predicted equations of motion for GR [28, 29]. The estimator is designed to cancel the effect of galaxy bias on linear scales. To sample the velocity field a redshift space distortion (RSD) analysis of the galaxy sample is used to translate the measured galaxy clustering to velocities using the RSD parameter . On linear scales and through matter conservation this parameter relates the divergence of the peculiar velocity field with the overdensities as . The RSD analysis requires accurate redshifts and therefore this estimator is especially applicable to spectroscopic galaxy samples like those from SDSS BOSS. The first measurement of the statistic used an estimator based on galaxy weak lensing and galaxy clustering measurements [30]. In Pullen et al. [28] an estimator for the statistic using CMB lensing and galaxy clustering was proposed. This was expanded upon and revised in Wenzl et al. [26] to increase the overall accuracy and model several key systematic effects. We will use this revised estimator to measure in this work.
@@
Previous analyses have investigated the statistic using CMB lensing information from Planck and the SDSS BOSS samples [31, 32, 26]. Some of these analyses claimed a significant tension with the CDM prediction [31] while some found the measurement to be consistent with CDM within statistical expectations [32, 26], motivating further investigations. By leveraging independent CMB lensing measurements based on ACT observations we can derive new constraints providing a powerful consistency check. Additionally, the two lensing datasets are highly complementary: the Planck lensing information covers a larger sky area while the ACT dataset has a significantly higher angular resolution. By analyzing the two datasets using the same pipeline, and after checking for consistency between them, we derive a combined constraint resulting in the highest signal-to-noise constraint on with CMB lensing and galaxy data to date. We will put our results in context with constraints on from the literature at higher redshift using SDSS quasars [33] and ones using galaxy weak lensing instead of CMB lensing [30, 34, 35, 36, 37, 32, 38]. Careful investigations of the currently available datasets represent important preparatory work to develop and validate analysis pipelines applicable to upcoming datasets which will allow significantly improved constraints [28]. These include the Simons Observatory (SO) [39] and CMB-S4 [40] for measurements of CMB lensing and for galaxy samples include the Dark Energy Spectroscopic Instrument (DESI) [41], Euclid [42], LSST [43], the Spectro-Photometer for the History of the Universe, Epoch of Reionization, and Ices Explorer (SPHEREx) [44] and the Roman Space Telescope [45].
@@
In Section II the key results of the analysis are summarized. The definitions of the angular power spectra and the statistic and estimator are summarized in Section III. In Section IV the datasets used and in Section V the blinding policy followed during development are described. The analysis pipeline for the cross-correlation measurement between ACT DR6 CMB lensing and the SDSS BOSS galaxy samples is described, validated, and applied in Section VI. In Section VII the measurements are used to derive new constraints on the statistic and in Section VIII a conclusion of the analysis is given.
II Summary of key results
In this work, we measure the angular cross-power spectrum, , between the ACT DR6 lensing map and the SDSS BOSS DR12 CMASS and LOWZ galaxy samples. Together with previously measured galaxy auto-correlations and RSD analyses for the RSD parameter , respectively constraining the growth and growth rate of structure, we report new constraints on the statistic. We compare the measured values with predictions for a fiducial cosmology based on the CDM fit using Planck 2018 CMB anisotropies and SDSS BOSS DR12 baryon acoustic oscillations (BAO) [2], in particular the reported value of .
We carried out the analysis following the ACT DR6 lensing blinding policy to minimize unconscious biases that may affect the results. Before the results were unblinded the analysis pipeline was validated and all analysis choices were finalized (see Section V for details) using an extensive set of consistency and null checks.
We perform an end-to-end test of our pipeline based on simulated map realizations, showing that the input spectra are accurately recovered. We show consistency between multiple distinct error estimation techniques using simulations and analytic calculations; we also check this consistency internally from the data themselves using a jackknife technique. We observe an approximately flat distribution of PTE (“probability to exceed”) values from our suite of tests indicating an accurate estimation of uncertainties. We account for magnification bias in our modeling throughout the analysis. For the high-resolution ACT map, the lensing reconstruction contamination from extragalactic foregrounds is a significant challenge as without mitigation the results would be impacted. Through several tests on simulations and null tests on the data, we show that the baseline ACT DR6 CMB lensing map, which uses a bias-hardened estimator [46, 47, 48, 20, 49], sufficiently mitigates the effect of bias from extragalactic foregrounds within the statistical constraining power of the cross-correlation analysis. We performed the full set of tests in the cosmological analysis range restricted to linear scales for the calculation. Additionally, we carried out the tests for an extended validation range including smaller scales that were not used in the estimation, making the measurement applicable to future analyses considering smaller scales. In addition to the ACT-only measurements we also present combined constraints with the Planck PR4 lensing information.
In Fig. 1 we show an overview of the measurements presented in this work. The results are shown as a function of scale; we also show combined constraints from all scales in the range used for the cosmological analysis, plotted at their effective redshift against the predictions for the fiducial cosmological model.
GR predicts the statistic to be independent as a function of angular scale in the linear regime (top panels of Fig. 1). Many alternative theories of gravity predict a scale dependence in the statistic (e.g. [27, 28]). We test all of our measurements for consistency with this prediction and find that all presented measurements are consistent with scale independence as predicted by GR within statistical expectations (). Just like for the measurements using Planck and SDSS CMASS [31, 32, 26], we find that the measurement tends low on large scales . However, the trend is not statistically significant given the statistical uncertainty even when combining the results.
Since there is no statistical evidence for scale dependence we can derive constraints combined across all multipole scales in the cosmological analysis range. We find
where the uncertainties show the confidence range and the subdominant systematic error budget is estimated to be 3% () and 4% () respectively. For each, is the effective redshift of the measurement. We find that our constraints on are statistically consistent with previous measurements presented in Wenzl et al. [26] using Planck PR4 lensing and SDSS BOSS at the same redshifts. Since they are consistent we combine the constraints considering the correlation between the two measurements. We find
giving the strongest constraints on with CMB lensing and galaxy clustering to date. These measurements are general for a wide range of cosmological models including GR and several extensions (see Section III). From constraints on the expansion history of the Universe, these gravitational models make direct predictions for the value of over time and can be compared with these measurements.
We test these measurements for statistical consistency with GR predictions based on our CDM fiducial cosmology. The predicted values are given as
where the uncertainty is dominated by the constraint on for our fiducial cosmology. We find that in all four cases, the measurements are statistically consistent with GR predictions ().
III Background
III.1 Angular power spectra
@@
where refers to clustering and refers to CMB lensing. is the matter power spectrum. Note here . Under the linear galaxy bias approximation used throughout this work, we have and . Each tracer we consider has a specific kernel function that characterizes the distances to which it is sensitive. The general lensing kernel for a source at redshift is given by
@@
We calculate the theoretical predictions for and for the fiducial cosmology in this work using CCL 111https://github.com/LSSTDESC/CCL [58] and the underlying 3D matter power spectrum using CAMB [59, 60].
III.2 Gravity Statistic EG
The gravity statistic tests the Poisson equation as predicted for GR by comparing the divergence of the peculiar velocity field, , with gravitational lensing . Under the assumption of isotropy, homogeneity, and flatness, the universe can be described on large scales with the perturbed Friedmann-Lemaître-Robertson-Walker (FLRW) metric . Here is the scale factor, with today, describes perturbations in the time component, and perturbations in the spatial component. The statistic is defined as [27]
@@ III.2 Gravity Statistic EG
where is the fractional matter density today compared to the critical density. In CDM specifically, the growth rate is, to high accuracy, directly given by [61], so that the statistic can be predicted from a measurement of . In GR more broadly the growth rate is generally scale-independent on linear scales which represents another testable prediction of the theory. Alternative descriptions of gravity like gravity [8] and Chameleon gravity [10], while matching the background expansion of the universe, predict deviations in the growth history and therefore the statistic. In general, both the expected value of changes and scale dependence can be introduced in these alternative frameworks (see e.g. [27, 28]).
@@
where is the effective redshift of the measurement, is the Hubble parameter at the effective redshift and is a correction for the effect of magnification bias that is described in detail in Section VI.1.5 and is based on the results of [57, 26]. The estimator combines the observed angular auto-power spectrum and from an RSD analysis of the galaxy sample with the angular cross-power spectrum with CMB lensing .
@@
for CMB lensing combined with the SDSS CMASS and LOWZ galaxy samples [26]. Astrophysical effects like galaxy bias evolution with redshift and magnification bias could limit the overall accuracy. Galaxy bias evolution with redshift is difficult to correct for since it is degenerate with the cosmological signal of interest. In Wenzl et al. [26] a CDM fit was used to constrain the size of this systematic bias and it was concluded that for CMB lensing combined with CMASS and LOWZ, the systematic error budget is 1% and 2% respectively. For the present analysis, this is well within the statistical uncertainty and we account for it in our overall systematic error budget. For future surveys with larger constraining power, narrower bins in redshift can be used to reduce the impact of this systematic. Another systematic effect is magnification bias which can be estimated based on the photometric selection of the galaxy samples and corrected for in the analysis. The effect partially cancels for as the cross- and auto-correlations are affected in the same direction. The effect is cumulative along the line of sight and therefore more relevant for higher redshift samples: for LOWZ the magnification bias is negligible, but for CMASS a 2–3% correction is needed which we discuss in detail in Section VI.1.5. The uncertainty of the correction is negligible compared to the other systematic error budgets and the statistical uncertainty [26].
IV Data
In this section, we summarize the observational datasets used in this work. In Section IV.1 we discuss the ACT CMB lensing map and auxiliary data products. In Section IV.2 the LOWZ and CMASS galaxy samples are described and we show how we build the galaxy overdensity maps for cosmological analysis. Finally, in Section IV.3 we describe the Planck CMB lensing map we use to obtain combined constraints from ACT and Planck + SDSS.
IV.1 CMB lensing from the Atacama Cosmology Telescope
For this analysis, we use the ACT DR6 CMB lensing map and auxiliary data products presented in Madhavacheril et al. [6]. In addition to the baseline lensing map, we use a range of data maps with alternative processing, a set of 400 simulated map realizations for each, the mask of the primary CMB observations, and the noise curve of the CMB lensing map. The alternative maps include CMB lensing reconstruction performed on the individual and maps as well as their difference using either temperature-only information or both temperature and polarization information. There is also a reconstruction that, instead of the baseline bias hardened estimator, uses CIB-deprojection as discussed in MacCrann et al. [49]. This alternative map has a slightly more restrictive mask which we account for when making comparisons to the baseline maps.
The simulations consist of 400 simulated CMB realizations with a known CMB lensing realization applied [63]. For each simulation, with the observational mask applied, the same reconstruction pipeline as used on the data was run which ensures that the simulations capture the full noise characteristics of the observed map. The same has been done for the alternative maps used for the analysis.
In MacCrann et al. [49] it was demonstrated, based on a range of tests including on the Websky CMB simulations [64, 65], that the estimator used for the ACT DR6 CMB lensing map sufficiently mitigates the contamination of foregrounds in the analysis for the angular auto-power spectrum. From this alone, however, it does not necessarily follow that there is no contamination for angular cross-power spectra with LSS datasets as calculated in this work. We performed extensive testing for foreground contamination in the analysis (see Sections VI.2.3, VI.2.4 and VI.2.5). Additionally, foreground contamination for the angular cross-power spectra with WISE galaxy samples was investigated in Farren et al. [63] where no significant effect was found.
We developed this analysis under ACT DR6 blinding guidelines for lensing analyses which means that throughout development the ACT DR6 lensing map used had an unknown factor of order applied that was only removed after finalizing the analysis choices as well as after investigating the consistency and null test results.
To ensure that large noise in our data on small scales does not affect our results we apply a low pass filter defined by using , filtering small scales far outside of the cosmological analysis range. We apply this filter to avoid signal bleed from small scales far outside the analysis range similar to comparable analyses [66, 63, 26]. We also restrict the CMB lensing map to regions where the apodized CMB mask is above 0.6 to avoid dividing the maps by small values.
IV.2 Galaxy catalog from SDSS
@@
In order to create simulated galaxy map realizations and to estimate the analytic covariance we need an estimate of the shot noise in the data given as [67, 25, 26]:
@@
IV.3 CMB lensing measurement from Planck

### v2

@@ I Introduction
While the current standard model of cosmology, CDM, fits a wide range of observational constraints (e.g. [1, 2, 3, 4, 5, 6]), the explanation behind the observed magnitude of the cosmological constant, , remains unresolved [7]. This has motivated the proposal of several alternative explanations, including a novel type of energy, called dark energy, and gravitational models including gravity [8], DGP [9], Chameleon gravity [10], TeVeS [11] and others (for a review see Clifton et al. [12]). These modified gravitational models are generally designed to match the observed homogeneous accelerated expansion of the late-time universe and employ screening mechanisms so that they match general relativity (GR) on solar system scales [13]. However, they do typically predict deviations from GR in the equations of motion affecting the growth of density fluctuations that seed large-scale structure (LSS) in the Universe [14]. Probing the growth of large-scale structure in the Universe, therefore, offers a powerful way to test gravity, allowing the construction and measurement of discriminating tests between CDM and alternative models of gravity (see e.g. [15, 13, 16, 17]).
@@
Photons from the Cosmic Microwave Background (CMB) get gravitationally lensed along the line of sight through the intervening gravitational potential formed from the large-scale structure of the Universe [18]. This CMB lensing effect can be detected in the CMB anisotropies through its characteristic effect of coupling different scales which would otherwise be independent from each other. CMB lensing maps have been created for recent CMB measurements from the Planck satellite [19], the Atacama Cosmology Telescope (ACT) [20, 21] and the South Pole Telescope (SPT) [22, 23]. The newest ACT CMB lensing map constructed based on the Data Release 6 (DR6) CMB dataset covers a total of and offers significantly higher resolution and signal-to-noise than previous maps based on Planck data [21].
@@
The cosmic volume probed by CMB lensing along the line of sight overlaps with the volumes probed by galaxy surveys. This allows measurements of the correlation between CMB lensing and galaxy clustering which serves as an opportunity to break degeneracies between the amplitude of fluctuations and the galaxy bias for galaxy clustering alone. The Sloan Digital Sky Survey III Baryon Oscillation Spectroscopic Survey (SDSS BOSS) DR12 contains two large spectroscopic galaxy samples, CMASS and LOWZ which may be used to study this correlation, the release used is the latest and final version of the data [24, 25, 1]. The correlation has been previously measured with the smaller ACT DR4 CMB lensing map in Darwish et al. [20], Marques et al. [26]. Combining the SDSS BOSS galaxy catalogs with the ACT DR6 CMB lensing map offers the opportunity to derive new competitive constraints on the correlation between CMB lensing and galaxy clustering, independent of Planck, and allows new constraints on gravity. To demonstrate the accuracy of this correlation we perform a detailed suite of null and consistency tests to show that biases from systematic effects like extragalactic foregrounds and magnification bias are mitigated and accurately accounted for. For this work, we focus on using these cross-correlation measurements for constraints on gravity and make them available for further analysis.
@@
The measurement of this cross-correlation, combined with previous analyses of the galaxy data in Wenzl et al. [27], allow us to constrain the so-called statistic proposed in Zhang et al. [28]. This statistic builds a discriminating test of gravity by comparing the divergence of the peculiar velocity field with a measurement of gravitational lensing . The proposed ratio compares the two sides of the generalized Poisson equation and tests the equivalence of the spatial and temporal gravitation potentials, directly testing the predicted equations of motion for GR [29, 30]. The estimator is designed to cancel the effect of galaxy bias on linear scales. To sample the velocity field a redshift space distortion (RSD) analysis of the galaxy sample is used to translate the measured galaxy clustering to velocities using the RSD parameter . On linear scales and through matter conservation this parameter relates the divergence of the peculiar velocity field with the overdensities of galaxies as . The RSD analysis requires accurate redshifts and therefore this estimator is especially applicable to spectroscopic galaxy samples like those from SDSS BOSS. The first measurement of the statistic used an estimator based on galaxy weak lensing and galaxy clustering measurements [31]. In Pullen et al. [29] an estimator for the statistic using CMB lensing and galaxy clustering was proposed. This was expanded upon and revised in Wenzl et al. [27] to increase the overall accuracy and model several key systematic effects. We will use this revised estimator to measure in this work.
@@
Previous analyses have investigated the statistic using CMB lensing information from Planck and the SDSS BOSS samples [32, 33, 27]. Some of these analyses claimed a significant tension with the CDM prediction [32] while some found the measurement to be consistent with CDM within statistical expectations [33, 27], motivating further investigations. By leveraging independent CMB lensing measurements based on ACT observations we can derive new constraints providing a powerful consistency check. Additionally, the two lensing datasets are highly complementary: the Planck lensing information covers a larger sky area while the ACT dataset has a significantly higher angular resolution. By analyzing the two datasets using the same pipeline, and after checking for consistency between them, we derive a combined constraint resulting in the highest signal-to-noise constraint on with CMB lensing and galaxy data to date. We will put our results in context with constraints on from the literature at higher redshift using SDSS quasars [34] and ones using galaxy weak lensing instead of CMB lensing [31, 35, 36, 37, 38, 33, 39]. Careful investigations of the currently available datasets represent important preparatory work to develop and validate analysis pipelines applicable to upcoming datasets which will allow significantly improved constraints [29]. These include the Simons Observatory (SO) [40] and CMB-S4 [41] for measurements of CMB lensing and for future spectroscopic samples such as from the Dark Energy Spectroscopic Instrument (DESI) [42] and Euclid.
@@
In Section II we summarize the background for this work, giving definitions of the angular power spectra and the statistic and estimator used in this work. In Section III we describe the datasets used and in Section IV the blinding policy followed during development. In Section V we describe how the analysis pipeline for the cross-correlation measurement between ACT DR6 CMB lensing and the SDSS BOSS galaxy samples is validated and applied. In Section VI we use the measurements to derive new constraints on the statistic and in Section VII we give a conclusion of the analysis.
II Background
II.1 Angular power spectra
@@
where refers to clustering and refers to CMB lensing. Note here . Under the linear galaxy bias approximation used throughout this work, we have and where is the matter power spectrum. Each tracer we consider has a specific kernel function that characterizes the distances to which it is sensitive. The general lensing kernel for a source at redshift is given by
@@
We calculate the theoretical predictions for and for the fiducial cosmology in this work using CCL 111https://github.com/LSSTDESC/CCL [51] and the underlying nonlinear 3D matter power spectrum using CAMB [18, 52].
II.2 Gravity Statistic EG
The gravity statistic tests the Poisson equation as predicted for GR by comparing the divergence of the peculiar velocity field, , with gravitational lensing . Under the assumption of isotropy, homogeneity, and flatness, the universe can be described on large scales with the perturbed Friedmann-Lemaître-Robertson-Walker (FLRW) metric . Here is the scale factor, with today, describes perturbations in the time component, and perturbations in the spatial component [53]. The statistic is defined as [28]
@@ II.2 Gravity Statistic EG
where is the fractional matter density today compared to the critical density. In CDM specifically, the growth rate is, to high accuracy, directly given by [54], so that the statistic can be predicted from a measurement of . In GR more broadly the growth rate is generally scale-independent on linear scales which represents another testable prediction of the theory. Alternative descriptions of gravity like gravity [8, 55] and Chameleon gravity [10], while matching the background expansion of the universe, predict deviations in the growth history and therefore the statistic. In general, both the expected value of changes and scale dependence can be introduced in these alternative frameworks (see e.g. [28, 29]).
@@
where is the effective redshift of the measurement, is the Hubble parameter at the effective redshift and is a correction for the effect of magnification bias that is described in detail in Section V.1.4 and is based on the results of [50, 27]. The estimator combines the observed angular auto-power spectrum and from an RSD analysis of the galaxy sample with the angular cross-power spectrum with CMB lensing .
@@
for CMB lensing combined with the SDSS CMASS and LOWZ galaxy samples [27]. Astrophysical effects like galaxy bias evolution with redshift and magnification bias could limit the overall accuracy. Galaxy bias evolution with redshift is difficult to correct for since it is degenerate with the cosmological signal of interest. In Wenzl et al. [27] a CDM fit was used to constrain the size of this systematic bias and it was concluded that for CMB lensing combined with CMASS and LOWZ, the systematic error budget is 1% and 2% respectively. For the present analysis, this is well within the statistical uncertainty and we account for it in our overall systematic error budget. For future surveys with larger constraining power, narrower bins in redshift can be used to reduce the impact of this systematic. Another systematic effect is magnification bias which can be estimated based on the photometric selection of the galaxy samples and corrected for in the analysis. The effect partially cancels for as the cross- and auto-correlations are affected in the same direction. The effect is cumulative along the line of sight and therefore more relevant for higher redshift samples: for LOWZ the magnification bias is negligible, but for CMASS a 2–3% correction is needed which we discuss in detail in Section V.1.4. The uncertainty of the correction is negligible compared to the other systematic error budgets and the statistical uncertainty [27].
III Data
III.1 Galaxy catalog from SDSS
@@
To create simulated galaxy map realizations and to estimate the analytic covariance we need an estimate of the shot noise in the data given as [56, 26, 27]:
@@
III.2 CMB lensing from the Atacama Cosmology Telescope
For this analysis, we use the ACT DR6 CMB lensing map and auxiliary data products presented in Madhavacheril et al. [6]. In addition to the baseline lensing map, we use a range of data maps with alternative processing, the mask of the primary CMB observations, and the noise curve of the CMB lensing map. The alternative maps include CMB lensing reconstruction performed on the individual and maps as well as their difference using either temperature-only information or both temperature and polarization information. There is also a reconstruction that, instead of the baseline bias hardened estimator, uses CIB-deprojection as discussed in MacCrann et al. [58]. This alternative map has a slightly more restrictive mask which we account for when making comparisons to the baseline maps.
In MacCrann et al. [58] it was demonstrated, based on a range of tests including on the Websky CMB simulations [62, 63], that the estimator used for the ACT DR6 CMB lensing map sufficiently mitigates the contamination of foregrounds in the analysis for the angular auto-power spectrum. From this alone, however, it does not necessarily follow that there is no contamination for angular cross-power spectra with LSS datasets as calculated in this work. We performed extensive testing for foreground contamination in the analysis (see Sections V.2.3, V.2.4 and V.2.5). Additionally, foreground contamination for the angular cross-power spectra with WISE galaxy samples was investigated in Farren et al. [64] where no significant effect was found.
We apply a low pass filter defined by using , to avoid signal bleed from small scales far outside the analysis range similar to comparable analyses [65, 64, 27]. We also restrict the CMB lensing map to regions where the apodized CMB mask is above 0.6 to avoid dividing the maps by small values.
III.3 Simulated map realizations
The ACT DR6 lensing data products include 400 simulated CMB lensing map realizations [6]. The simulations consist of 400 simulated CMB realizations with a known CMB lensing realization applied [64]. For each simulation, with the observational mask applied, the same reconstruction pipeline as used on the data was run which ensures that the simulations capture the full noise characteristics of the observed map. This makes them ideal for testing our analysis pipeline. The same has been done for the alternative maps used for the analysis.
Using the input realizations that were used to create the simulated ACT maps we create 400 simulated Gaussian galaxy overdensity map realizations for both CMASS and LOWZ that have the expected correlation with the simulated CMB lensing maps, the expected auto-correlation, and the same noise level as the observed galaxies given by the shot noise estimate. As done for the data we create two versions of the galaxy maps for the auto- and cross-correlation where for the cross-correlation we account for the shifted from the additional reweighting we do for the cross-correlation only [Eq. 18]. This approach is described in detail in Wenzl et al. [27], building on previous work [66, 64]. We use linear galaxy biases of for CMASS and for LOWZ which were inferred by comparing the measured angular auto-power spectra of the galaxies to the amplitude of the input theory angular power spectra calculated for our fiducial cosmology. Before unblinding the data we used a preliminary value of [67, 68]. We use the simulated maps to estimate the baseline covariance for the angular power spectra and show that our analysis accurately recovers the input angular-power spectra.
III.4 CMB lensing measurement from Planck